# M3-VOS: Multi-Phase, Multi-Transition, and Multi-Scenery Video Object Segmentation

### v1

@@ Abstract
Intelligent robots need to interact with diverse objects across various environments. The appearance and state of objects frequently undergo complex transformations depending on the object properties, e.g., phase transitions. However, in the vision community, segmenting dynamic objects with phase transitions is overlooked. In light of this, we introduce the concept of phase in segmentation, which categorizes real-world objects based on their visual characteristics and potential morphological and appearance changes. Then, we present a new benchmark, Multi-Phase, Multi-Transition, and Multi-Scenery Video Object Segmentation (M3-VOS), to verify the ability of models to understand object phases, which consists of 479 high-resolution videos spanning over 10 distinct everyday scenarios. It provides dense instance mask annotations that capture both object phases and their transitions. We evaluate state-of-the-art methods on M3-VOS, yielding several key insights. Notably, current appearance-based approaches show significant room for improvement when handling objects with phase transitions. The inherent changes in disorder suggest that the predictive performance of the forward entropy-increasing process can be improved through a reverse entropy-reducing process. These findings lead us to propose ReVOS, a new plug-and-play model that improves its performance by reversal refinement. Our data and code will be publicly available.
@@ 1 Introduction
Given the data, we analyze the current understanding of object state changes in the field of computer vision from the perspective of Video Object Segmentation (VOS). On one hand, VOS task requires the visual model to possess a comprehensive understanding of object phase knowledge to track objects that undergo significant changes in appearance and morphology during phase transitions. On the other hand, video object segmentation serves as an upstream task for other tasks, such as recognizing higher-level attributes and affordances and object manipulation tasks.
Equipped with our benchmark, we analyze state-of-the-art VOS algorithms [2, 4, 33, 31, 24]. We found significant room for improvement in the understanding of objects undergoing phase transitions. Additionally, by analyzing the performance differences across cases, we identified potential challenges in phase changes and proposed specific improvements, establishing a baseline ReVOS utilizing reverse memory, booster, and readout fusion to advance object segmentation with phase transformations.
We conclude in Sec. 6.2 by analyzing the different performances of the existing models facing different categories of phase transitions. We hope this work will motivate further exploration into more robust object knowledge understanding. In summary, our main contributions are as follows:
@@ 2.2 Video Object Segmentation
Video object segmentation is the process of segmenting target foreground objects from the video background at the pixel level within a video sequence. The target foreground object is typically specified either by providing a mask in the first frame, known as semi-supervised VOS [20, 29, 22]. This task requires the effective tracking of the target and a clear understanding of its boundaries, which demands that models have a thorough comprehension of the object’s appearance [20, 29], dynamic properties [28, 32], and motion information [7] within the video, as well as other physical characteristics. This is precisely why we chose video object segmentation as the specific task to evaluate the ability of vision models to understand object phase transitions.
@@ 4.2 Annotation Design
In the annotation, we instructed volunteers to use the information provided by the video collectors about the target objects and the semi-auto Annotation Tool (detailed in Sec. 4.3.1) to annotate the masks of the target objects in each frame of the video. For various uncertain void masks caused by unclear image quality or motion blur, we instructed volunteers to annotate the corresponding void masks, similar to the approach in the VOST [28]. In this way, we can disregard the influence of these pixels during the evaluation process. Volunteers were also required to select the initial and final phases of the target objects from Tab. 2 based on their states in the video, and subsequently label the corresponding phase transformations.
@@ 4.3 Annotation Collection
In this section, we introduce our semi-auto Annotation tool which efficiently annotates the masks of target objects in the videos and our annotation collection pipeline.
@@ 4.3.1 Semi-Auto Multi-Level Annotation Tool
Our annotation tool mainly employs a paint-and-erase approach. Volunteers can use a brush and eraser tool, along with a magnification feature, to accurately annotate the masks of target objects and the void masks in the video. To achieve high-frame-rate mask annotation, we integrated the fully automated annotation tool provided by Cutie [4] based on RITM [26] for interactive image segmentation and Cutie for video object segmentation. After volunteers annotate masks for a specific frame using the paint-and-erase approach or point prompt method, they can utilize Cutie to propagate masks for that frame. Subsequently, the propagated pseudo-ground truth can refined using the paint-and-erase approach. The neural network based automatic annotation tool provides good pseudo ground truth references for objects with clear contours. However, the tool often fails to provide satisfactory pseudo-ground truth for target objects like water vapor, thick smoke, and splashes, which have unclear outlines. Thus, our tool integrates a color difference masking, allowing users to select a specific pixel value in the image and set a range to create a color difference template for the video. Given a pixel $p_{s}$ and any other pixel $p_{i}$, the position of $p_{i}$ will be labeled according to
@@
We label the pixel $p_{i}$ if and only if the color difference between $p_{s}$ and $p_{i}$ is within $0.1\delta$, and the differences in saturation and brightness are within $\delta$. Here, $\delta$ is a user-adjustable parameter. Our testing has shown that this functionality effectively annotates objects with unclear contours but distinct colors, especially for gases and liquids.
@@ 4.5 Core Subset
To facilitate analysis of the challenge facing the phase transition in the VOS task, we extracted video clips that humans consider to be highly representative and comprehensive in variety. Finally, we split M3-VOS into two subsets for evaluation, including full subset and core subset. The details can be found in Supplementary.
@@ 12.1 IoU Analysis
In terms of model selection of the dual-model cross-validation process, we adapt the annotation model to the latest SAM2 [24]. We utilized the open-source base plus model configuration and checkpoints, as this configuration is more effective in fully segmenting our target objects compared to other model setups.
@@ 13 Multi-Level Semi-Auto Annotation Tool
In Fig. 9, we show our details of the interactive UI of the multi-level semi-auto annotate tool. We implement this tool based on the interactive demo from Cutie [4], including pixel level, appearance-level, and object level. In particular, we implement the object-level function using the SAM2 model and Cutie model. In this way, we could perform the dual-model cross-validate analysis in Sec. 12.
@@ 14 Details of Core Subset
We extract a subset of cases that better represent the full dataset and refer to it as core subset. For each specific scenario, we extracted a subset of cases. During selection of core subset in each scenario, we consider a series of factors: number of the full set, number of classes included, the difficulty of cases. As is shown in Tab. 10, We choose size of core subset of each scenario to make it closer to the proportion of the full set and the class number.
@@ 15.3 Relation between Challenge and Performance
Similarly, the curve in Fig. 10(b) indicates that the relative velocity shows a positive correlation with segmentation difficulty. We observe that when the velocity is more extreme, either too slow or too fast, the performance improvement of ReVOS-Cutie becomes more significant.
@@ 16 More Failure Cases
In case 2 (assemble puzzles) of Fig. 11, the intra-solid transition with multi-instances usually instance confusion because of the similarity distribution. However, our methods is more robust when facing the intra-solid transition with multi-instances.
In case 1 (cut apples) of Fig. 12, the tracking loss and mask incompleteness usually happen when the white pulp leaks out. The performance in the intra-solid phase transition with the color change challenge is not so well.
In case 2 (pour tea) of Fig. 12, the flow of tea liquid into the tea cup is always accompanied by tracking loss. Besides, transparent teapots and tea liquids also suffer from similar interference and are confused. Although our method improves this situation slightly, this multi-challenge case still has improved space.

### v3

@@ Abstract
Intelligent robots need to interact with diverse objects across various environments. The appearance and state of objects frequently undergo complex transformations depending on the object properties, e.g., phase transitions. However, in the vision community, segmenting dynamic objects with phase transitions is overlooked. In light of this, we introduce the concept of phase in segmentation, which categorizes real-world objects based on their visual characteristics and potential morphological and appearance changes. Then, we present a new benchmark, Multi-Phase, Multi-Transition, and Multi-Scenery Video Object Segmentation (M3-VOS), to verify the ability of models to understand object phases, which consists of 479 high-resolution videos spanning over 10 distinct everyday scenarios. It provides dense instance mask annotations that capture both object phases and their transitions. We evaluate state-of-the-art methods on M3-VOS, yielding several key insights. Notably, current appearance-based approaches show significant room for improvement when handling objects with phase transitions. The inherent changes in disorder suggest that the predictive performance of the forward entropy-increasing process can be improved through a reverse entropy-reducing process. These findings lead us to propose ReVOS, a new plug-and-play model that improves its performance by reversal refinement. Our data and code will be publicly available at https://zixuan-chen.github.io/M-cube-VOS.github.io/.
@@ 1 Introduction
Given the data, we analyze the current understanding of object state changes in the field of computer vision from the perspective of Video Object Segmentation (VOS). On one hand, VOS task
requires the visual model to possess a comprehensive understanding of object phase knowledge to track objects that undergo significant changes in appearance and morphology during phase transitions. On the other hand, video object segmentation serves as an upstream task for other tasks, such as recognizing higher-level attributes and affordances and object manipulation tasks.
Equipped with our benchmark, we analyze state-of-the-art VOS algorithms [2, 4, 35, 33, 25]. We found significant room for improvement in the understanding of objects undergoing phase transitions.
Additionally, by analyzing the performance differences across cases, we identified potential challenges in phase changes and proposed specific improvements, establishing a baseline ReVOS utilizing reverse memory, booster, and readout fusion to advance object segmentation with phase transformations.
We conclude in Sec. 6.2 by analyzing the different performances of the existing models facing different categories of phase transitions. We hope this work will motivate further exploration into more robust object knowledge understanding.
In summary, our main contributions are as follows:
@@ 2.2 Video Object Segmentation
Video object segmentation is the process of segmenting target foreground objects from the video background at the pixel level within a video sequence. The target foreground object is typically specified either by providing a mask in the first frame, known as semi-supervised VOS [21, 30, 23].
This task requires the effective tracking of the target and a clear understanding of its boundaries, which demands that models have a thorough comprehension of the object’s appearance [21, 30], dynamic properties [29, 34], and motion information [7] within the video, as well as other physical characteristics. This is precisely why we chose video object segmentation as the specific task to evaluate the ability of vision models to understand object phase transitions.
@@ 4.2 Annotation Design
In the annotation, we instructed volunteers to use the information provided by the video collectors about the target objects and the semi-auto Annotation Tool (detailed in Sec. 4.3.1) to annotate the masks of the target objects in each frame of the video. For various uncertain void masks caused by unclear image quality or motion blur,
we instructed volunteers to annotate the corresponding void masks, similar to the approach in the VOST [29]. In this way, we can disregard the influence of these pixels during the evaluation process. Volunteers were also required to select the initial and final phases of the target objects from Tab. 2 based on their states in the video, and subsequently label the corresponding phase transformations.
@@ 4.3 Annotation Collection
In this section, we introduce our semi-auto Annotation tool which efficiently annotates the masks of target objects in the videos, and our annotation collection pipeline.
@@ 4.3.1 Semi-Auto Multi-Level Annotation Tool
Our annotation tool mainly employs a paint-and-erase approach. Volunteers can use a brush and eraser tool, along with a magnification feature, to accurately annotate the masks of target objects and the void masks in the video. To achieve high-frame-rate mask annotation, we integrated the fully automated annotation tool provided by Cutie [4] based on RITM [27] for interactive image segmentation and Cutie for video object segmentation. After volunteers annotate masks for a specific frame using the paint-and-erase approach or point prompt method, they can utilize Cutie to propagate masks for that frame. Subsequently, the propagated pseudo-ground truth can refined using the paint-and-erase approach. The neural network based automatic annotation tool provides good pseudo ground truth references for objects with clear contours. However, the tool often fails to provide satisfactory pseudo-ground truth for target objects like water vapor, thick smoke, and splashes, which have unclear outlines. Thus, our tool integrates a color difference masking, allowing users to select a specific pixel value in the image and set a range to create a color difference template for the video.
Given a pixel $p_{s}$ and any other pixel $p_{i}$, the position of $p_{i}$ will be labeled according to
@@
We label the pixel $p_{i}$ if and only if the color difference between $p_{s}$ and $p_{i}$ is within $0.1\delta$, and the differences in saturation and brightness are within $\delta$. Here, $\delta$ is a user-adjustable parameter.
Our testing has shown that this functionality effectively annotates objects with unclear contours but distinct colors, especially for gases and liquids.
@@ 4.5 Core Subset
To facilitate analysis of the challenge facing the phase transition in the VOS task, we extracted video clips that humans consider to be highly representative and comprehensive in variety. Finally, we split M3-VOS into two subsets for evaluation, including the full subset and core subset. The details can be found in Supplementary.
@@ 12.1 IoU Analysis
In terms of model selection of the dual-model cross-validation process, we adapt the annotation model to the latest SAM2 [25]. We utilized the open-source base plus model configuration and checkpoints, as this configuration is more effective in fully segmentation [25, 9, 32] of our target objects compared to other model setups.
@@ 13 Multi-Level Semi-Auto Annotation Tool
In Fig. 9, we show our details of the interactive UI of the multi-level semi-auto annotate tool. We implement this tool based on the interactive demo from Cutie [4], including pixel level, appearance level, and object level. In particular, we implement the object-level function using the SAM2 model and Cutie model. In this way, we could perform the dual-model cross-validate analysis in Sec. 12.
@@ 14 Details of Core Subset
We extract a subset of cases that better represent the full dataset and refer to it as a core subset. For each specific scenario, we extracted a subset of cases. During the selection of the core subset in each scenario, we consider a series of factors: the number of the full set, the number of classes included, and the difficulty of cases. As is shown in Tab. 10, We choose the size of the core subset of each scenario to make it closer to the proportion of the full set and the class number.
@@ 15.3 Relation between Challenge and Performance
Similarly, the curve in Fig. 10(b) indicates that the relative velocity positively correlates with segmentation difficulty. We observe that when the velocity is more extreme, either too slow or too fast, the performance improvement of ReVOS-Cutie becomes more significant.
@@ 16 More Failure Cases
In case 2 (assemble puzzles) of Fig. 11, the intra-solid transition with multi-instances usually instance confusion because of the similarity distribution. However, our methods are more robust when facing the intra-solid transition with multi-instances.
In case 1 (cut apples) of Fig. 12, the tracking loss and mask incompleteness usually happen when the white pulp leaks out. The performance in the intra-solid phase transition with the color change challenge is not so good.
In case 2 (pour tea) of Fig. 12, the flow of tea liquid into the tea cup is always accompanied by tracking loss. Besides, transparent teapots and tea liquids are confused and suffer from similar interference. Although our method improves this situation slightly, this multi-challenge case still has improved space.